OpenStudy (keana):

Solve 2x^2 = -10x. A:x = 2, x = -5 B:x = 2, x = 5 C:x = 0, x = 5 D:x = 0, x = -5

OpenStudy (lgbasallote):

put 10x on the left side first \[2x^2 + 10x = 0\] now facotr! \[2x(x + 10) = 0\] now equate each factor to zero and solve for x solve for x when 2x = 0 solve for x when x + 10 = 0

OpenStudy (lgbasallote):

do you get it?

mathslover (mathslover):

\[2x^2=-10x\] \[2x^2/(-10)=x\] \[-x^2/5=x\] \[-x^2=5x\] \[-x^2/x=5\] \[-x=5\] \[x=-5\]

OpenStudy (lgbasallote):

careful here...you shouldn't divide by x because you will be missing one factor ;)